OpenStudy (anonymous):

equivalent to -3(2x-1)=-39

OpenStudy (bahrom7893):

-6x = -36 x=6

OpenStudy (anonymous):

x=7

OpenStudy (bahrom7893):

lol mad i was off by a sign, good job!

OpenStudy (anonymous):

^.^ you don't realize how happy I am lolololol

OpenStudy (anonymous):

Yeah, change the signs when you switch sides.

OpenStudy (saifoo.khan):

\[2x-1 = \frac{-39}{-3}\] \[2x - 1 = 13\] \[2x = 13 +1\] \[2x = 14\] \[x = \frac{14}{2}\] x = Seven

OpenStudy (anonymous):

thanks